The lunar surface reaches 370 K by midday. The atmosphere consists of neon, argon and traces of helium at a total pressure of only 2 x 10-14 atm. Calculate the root mean square speed of each component in the lunar atmosphere


1
Expert's answer
2021-04-07T11:30:11-0400

T = 370 K

MrNe = 20.18 g/mol = 0.02018 kg/mol

MrAr = 39.95 g/mol = 0.03995 kg/mol

MrHe = 4 g/mol = 0.004 kg/mol

R = 8.314

Vrms = "\sqrt{\frac{3RT}{M}}"

VrmsNe = "\sqrt{\frac{3\times 8.314\times370}{0.02018}}" = 676.24 m/s

VrmsAr = "\sqrt{\frac{3\times 8.314\times370}{0.03995}}" = 480.63 m/s

VrmsHe = "\sqrt{\frac{3\times 8.314\times370}{0.004}}" = 1518.93 m/s